# Bounds on $k$-hash distances and rates of linear codes

### v2

@@ II-B Upper Bound
In this section, we present a re-derivation of (7) and (8). The main tool used both in [18] and [19] is essentially some form of Jamison’s bound [22] (Lemma 1 below), which is also the same tool we use. The main difference is that, by applying an idea first introduced in [23], we obtain a shorter and simpler proof that directly relate the rate of the code to its minimum distance. This simpler approach is the starting point for the extension to the general case $q\geq k\geq 3$, which is then presented in the next section.
@@
Let $C$ be a linear trifferent code of dimension $m$ and length $n$. Let $G$ be the $m\times n$ generator matrix, let $d$ be the minimum Hamming distance of the code, and let $x$ be a codeword of weight $d$. Finally call $u\in{\mathbb{F}}_{3}^{m}$ the information vector associated to $x$, that is, assume $x=uG$, and also assume without loss of generality (by appropriate sorting and re-scaling of the columns of $G$) that $x$ has 00s in the last $n-d$ coordinates and $1$s in the first $d$ coordinates. Then, since the code is trifferent, any codeword different from 00 and $x$ must have a coordinate equal to $2$ among the first $d$ ones. So, if we call $g_{i}$ the $i$-th column of $G$, the $d$ affine subspaces defined by
$H_{i}=\{v\in{\mathbb{F}}_{3}^{m}:v\cdot g_{i}=2\}\,,i=1,\ldots,d$
cover the set ${\mathbb{F}}_{3}^{m}$ with the exception of 00 and $u$. Adding another subspace $H_{d+1}=\{v\in{\mathbb{F}}_{3}^{m}:v\cdot g_{1}=1\}$ we also cover $u$, still leaving out 00. By Jamison’s bound, $d+1\geq 2m$. In terms of rates and relative minimum distance $\delta=d/n$ this becomes
$R\leq\frac{1}{2}\delta+o(1)\,.$ (10)
@@
$\delta\leq\frac{2}{3}(1-R)+o(1)$
gives $R\leq(1-R)/3+o(1)$, which is asymptotically $R\leq 1/4$, essentially equivalent to [^3] Strictly speaking, to obtain the positive $\epsilon$ in (7) we need the fact that the Plotkin bound is not tight at positive rates. (7). The stronger bound (8) is obtained instead by using the best known bound on $\delta$, which is the linear programming bound of [24] adapted to $q$-ary codes [17], defined implicitly in $\delta$ by the inequality $R\leq R_{LP1}(q,\delta)$ with $q=3$ and
@@
$H_{q}(t)=t\log_{q}(q-1)-t\log_{q}t-(1-t)\log_{q}(1-t)\,.$ (12)
The bound on $R$ is found by combining (10) with $R\leq R_{LP}(3,\delta)$, which means solving $R=R_{LP1}(3,R/2)$.
@@ III General Case $q\geq k\geq 3$
We now turn to the case of $q$-ary linear $k$-hash codes proving our main result, which is a relation between the rate and minimum $k$-hash distances of codes. This will then be instantiated both to the case of asymptotic relative distance $\delta_{k}>0$ and to the case where only the requirement $d_{k}\geq 1$ is considered. To do this, we need the following generalization of Jamison’s bound to multiple coverings, due to Bruen [25].
Lemma 2 ([25])
Let $\mathcal{H}$ be a multiset of hyperplanes in ${\mathbb{F}}_{q}^{m}$. If no hyperplane in $\mathcal{H}$ contains 00 and each point in ${\mathbb{F}}_{q}^{m}\setminus\{0\}$ is covered by at least $t>0$ hyperplanes in $\mathcal{H}$, then
$|\mathcal{H}|\geq(m+t-1)(q-1)\,.$
Our application of this lemma to $k$-hashing is best presented in terms of the following basic result.
@@
Let $0,x_{1},x_{2},\ldots,x_{s-1}$ be $s$ codewords with $s$-hash distance $d_{s}>0$. Assume without loss of generality that they are all pairwise distinct in the first $d_{s}$ coordinates. Let $C^{\prime}$ be a subcode of dimension $(m-s+1)$ which only intersects trivially (in the origin) the subspace spanned by $x_{1},x_{2},\ldots,x_{s-1}$. Define, for $i=1,\ldots,d_{s}$, the set $S_{i}={\mathbb{F}}_{q}\setminus\{0,x_{1,i},x_{2,i},\ldots,x_{s-1,i}\}$. Let $G$ be the (full rank) $(m-s+1)\times n$ generator matrix of the subcode $C^{\prime}$ and $g_{i}$ be its $i$-th column. Consider the hyperplanes
$H_{i,b}=\{v\in{\mathbb{F}}_{q}^{m-s+1}\mid v\cdot g_{i}=b\}\,,\quad i=1,\ldots,d_{s}\,,\ b\in S_{i}\,.$
Since $|S_{i}|=q-s$, these are $(q-s)d_{s}$ hyperplanes, none of which contains the zero vector. Assume $d_{s+1}>0$; then, each $v\in{\mathbb{F}}_{q}^{m-s+1}\setminus\{0\}$ is covered at least $d_{s+1}$ times by those hyperplanes. Therefore, from Lemma 2 we have
$(q-s)d_{s}\geq(m-s+1+d_{s+1}-1)(q-1)\,.$
This is equivalent to
$d_{s+1}\leq\frac{q-s}{q-1}d_{s}-m+s\,,$
whenever the right hand side is positive. Otherwise, the assumption $d_{s+1}>0$ is impossible, and hence $d_{s+1}=0$. ∎
Remark 1
Note that Lemma 3 is an improvement on the recursive bound given in Theorem 3 whenever $d_{s}\geq q-1$. For $q=3$ and $s=2$, with $d_{3}\geq 1$, this lemma does not give exactly the procedure used in the previous section. The reason is that in the asymptotic regime $n\to\infty$, with $d_{s+1}$ potentially proportional to $n$, it is better to consider the subcode $C^{\prime}$ rather than covering the entire code using additional hyperplanes, as we did for $q=k=3$.
The main results of this paper derive immediately from an iteration of Lemma 3, which gives
$d_{k}\leq\frac{(q-2)^{\underline{k-2}}}{(q-1)^{k-2}}\left(d_{2}-\sum_{i=1}^{k-2}(m-i-1)\frac{(q-1)^{i}}{(q-2)^{\underline{i}}}\right)^{+}\,,$
Note that the last equation is an improvement upon the one given in (6).
Whenever we assume $m/n\to R$ and $d_{k}/n\to\delta_{k}>0$ as $n\to\infty$, the previous equation gives the following result.
@@
Consider linear codes in ${\mathbb{F}}_{q}^{n}$ of rate $R$ and minimum relative $k$-th hash distance $\delta_{k}>0$ and minimum Hamming relative distance $\delta_{2}$. Then, as $n\to\infty$,
$R\leq\frac{1}{\sum_{i=1}^{k-2}\frac{(q-1)^{i}}{(q-2)^{\underline{i}}}}\left(\delta-\frac{(q-1)^{k-2}}{(q-2)^{\underline{k-2}}}\delta_{k}\right)+O\left(\frac{1}{n}\right)\,.$
If we instead simply assume $d_{k}\geq 1$ as $n\to\infty$, we obtain the following.
@@
Consider linear $k$-hash codes in ${\mathbb{F}}_{q}^{n}$ of rate $R=m/n$ and relative minimum Hamming distance $\delta$. Then, as $n\to\infty$,
$R\leq\frac{\delta}{\sum_{i=1}^{k-2}\frac{(q-1)^{i}}{(q-2)^{\underline{i}}}}+O\left(\frac{1}{n}\right)\,.$
It can be easily seen that the upper bound that we found on $R$ improves the one of [1] given in Theorem 4 for every $q$ and $k$.
IV New upper bounds
In this section, we provide, using the result obtained in Corollary 1, new upper bounds on the rate of linear $k$-hash codes in ${\mathbb{F}}_{q}^{n}$ for every $q\geq k\geq 3$.
Using the Plotkin bound for $q\geq 3$ we obtain the following corollary.
@@ IV New upper bounds
Now, using the upper bound on $\delta$ of equation (14) and the bound for linear $k$-hash codes given in Corollary 1 we obtain
$R\leq\frac{q-1}{q}\frac{(1-R)}{\sum_{i=1}^{k-2}\frac{(q-1)^{i}}{(q-2)^{\underline{i}}}}+o(1)\,.$
Therefore, rearranging the terms we obtain the statement of the corollary. ∎
@@
where $\delta^{*}$ is the unique root of the following equation in $x$
$\frac{x}{\sum_{i=1}^{k-2}\frac{(q-1)^{i}}{(q-2)^{\underline{i}}}}=\\ H_{q}\left(\frac{1}{q}\left(q-1-(q-2)x-2\sqrt{(q-1)x(1-x)}\right)\right)\,,$
where $0\leq x\leq\frac{q-1}{q}$ and $H_{q}$ is the function defined in equation (12).
In Table I, we compare the bounds provided in Corollaries 23 and the one given in (4) for $k$=3 and $q\in[3,64]$. It can be seen that the linear programming bound performs better for $q\leq 19$ while for $q\geq 23$ the Plotkin bound gives a better result.
@@
$3$ $1/4=0.25$ 0.2198 0.3691
$4$ $1/3=0.\overline{3}$ 0.3000 $1/2=0.5$
$5$ $3/8=0.375$ 0.3441 0.5694
$7$ $5/12=0.41\overline{6}$ 0.3928 0.6438
$8$ $3/7=0.\overline{428571}$ 0.4080 $2/3=0.\overline{6}$
$9$ $7/16=0.4375$ 0.4200 0.6846
@@
$13$ $11/24=0.458\overline{3}$ 0.4497 0.7298
$16$ $7/15=0.4\overline{6}$ 0.4628 $3/4=0.75$
$17$ $15/32=0.46875$ 0.4663 0.7554
$19$ $17/36=0.47\overline{2}$ 0.4721 0.7646
$23$ $21/44=0.477\overline{27}$ 0.4811 0.7790
$25$ $23/48=0.4791\overline{6}$ 0.4846 0.7847
$27$ $25/52=0.48\overline{076923}$ 0.4877 0.7897
$29$ $27/56=0.482\overline{142857}$ 0.4903 0.7942
$31$ $29/60=0.48\overline{3}$ 0.4927 0.7982
$32$ $15/31\approx 0.483871$ 0.4938 $4/5=0.8$
$37$ $35/72=0.486\overline{1}$ 0.4984 0.8081
$41$ $39/80=0.4875$ 0.5013 0.8134
@@
$64$ $31/63=0.\overline{492063}$ 0.5119 $5/6=0.8\overline{3}$
Figure 1: Comparison between upper and lower bounds for $q\geq 5$ and $k=4$.
Remark 2
We observe that one could use the second linear programming bound or the straight-line bounds given in [17, 26] to improve the results of Corollaries 2 and 3 for different values of $q$ and $k$. Here we avoid to show those improvements to keep a simpler presentation of our results.
For a fixed value of $k$ and $q\to\infty$, both bounds given in Corollaries 2 and 3 converge to $1/(k-1)$, which is the same upper bound on the rate of $(q,k)$-hash codes (not necessarily linear) that one can derive from equation (5). Corollary 2 approaches $1/(k-1)$ from below since the rhs of equation (13) is strictly increasing in $q$ and since $\sum_{i=1}^{k-2}\frac{(q-1)^{i}}{(q-2)^{\underline{i}}}\geq k-2$, while Corollary 3 does not, see for example Table I where for $k=3$ and $q=41,\ldots,64$ we have upper bounds that exceed $1/2$.
@@
since the function $\log_{x}(\alpha x)$ is increasing in $x$ for $x\geq 2$ and $10<\alpha<$ 1 and since $k\geq 4$ and $q\geq k^{2}$. Then, by Bernoulli’s inequality we have that
$\frac{1}{2}\left(\frac{q-k+2}{q}\right)^{k-2}\geq\frac{1}{2}\left(1-\frac{(k-2)^{2}}{q}\right)\,.$
@@
Since the lhs of (15) is less than $1/(k-1)$, in order to prove the statement of the theorem we just need to show that
$\frac{1}{k-1}\leq\frac{1}{2}\left(1-\frac{(k-2)^{2}}{q}\right)\,,$
but this inequality is satisfied for ${q\geq k^{2}}$ and $k\geq 4$. ∎
@@
We conjecture that Theorem 6 still holds also if we relax the hypothesis to $q\geq 2k-3$ and $k\geq 3$. This would imply that, for all the interesting values of $q$ and $k$ (since the asymptotic rate of linear $k$-hash codes in ${\mathbb{F}}_{q}^{n}$ for $q\leq 2k-4$ is zero), our bound provides the best result.
In support of our conjecture, Table I provides an instance for $k=3$ where for every $q\geq 3$ the bound of Corollary 2 improves the one of equation (4) and Figure 1 reports the comparison between our bounds and the best known bounds in the literature for $k=4$ and $q\geq 2k-3$. In addition, we have numerically verified the conjecture for every $k\in[3,100]$ and $q\geq 2k-3$.
The authors in [1], using classical random coding techniques, provide the following achievability bound on the rate of linear $k$-hash codes in ${\mathbb{F}}_{q}^{n}$ for $q\geq\binom{k}{2}$:
$R\geq\min\bigg{\{}-\frac{1}{k-1}\log_{q}\left(1-\frac{q^{\underline{k}}}{q^{k}}\right),\frac{1}{k-2}\left(1-\log_{q}\binom{k}{2}\right)\bigg{\}}+o(1)\,,$ (16)
where it can be seen that for $q$ sufficiently larger than $k$ the minimum of (16) is achieved by the first term. This implies that for such values of $q$ and $k$, the lower bound coincides with the one for general codes given in equation (3). However, there is still a large gap between upper and lower bounds.
In Figure 1 , we compare our upper bounds and the lower bound of equation (16) for $k=4$ and $q\geq 2k-3$. We note that both the upper bounds of Corollaries 2, 3 and the lower bound (16) are asymptotically equal to $1/(k-1)$ for a fixed value of $k$ as $q\to\infty$.
V A Related Open Problem in Zero-Error Capacity

### v3

@@ II-B Upper Bound
In this section, we start by presenting a re-derivation of (8) and (9), showing how our proof extends to $d_{3}>1$. The main tool used both in [17] and [18] is essentially some form of Jamison’s bound [24] (Lemma 1 below), which is also the same tool we use. The main difference is that, by extending an idea first introduced in [25], we obtain a shorter and simpler proof that relates the rate of a linear trifferent code to its minimum $d_{2}$ (Hamming) distance. Combination with usual bounds on the minimum Hamming distance of codes gives the desired result. Apart form simplicity, the advantage is that this procedure is easily extended to $d_{3}>1$ as well as to the general case $q\geq k\geq 3$, which is discussed in the next Section.
@@
Let $C$ be a linear trifferent code of dimension $m$ and length $n$. Let $d_{2}$ be the minimum Hamming distance of the code, and let $x$ be a codeword of weight $d_{2}$. Assume without loss of generality (by appropriate sorting and re-scaling of the coordinates) that $x$ has 0’s in the last ndn-d coordinates and $1$’s in the first $d_{2}$ coordinates. Consider a linear subcode $C^{\prime}$ of dimension m1m-1 which only intersects trivially (in the origin) with the subspace spanned by $x$, and let $G$ be its $(m-1)\times n$ generator matrix. Then, since $C$ is trifferent, any non-zero codeword in $C^{\prime}$ must have a coordinate equal to $2$ among the first $d_{2}$ ones, for otherwise it is not trifferenciated with 0 and $x$. So, if we call $g_{i}$ the $i$-th column of $G$, the $d_{2}$ affine subspaces defined by
$H_{i}=\{v\in{\mathbb{F}}_{3}^{m-1}:v\cdot g_{i}=2\}\,,i=1,\ldots,d_{2}$ (17)
cover the set ${\mathbb{F}}_{3}^{m-1}$ with the exception of 0. By Jamison’s bound, $d_{2}\geq 2(m-1)$. In terms of rates and relative minimum distance $\delta_{2}=d_{2}/n$ this becomes
$R\leq\frac{1}{2}\delta_{2}+o(1)\,.$ (18)
@@
$\delta_{2}\leq\frac{2}{3}(1-R)+o(1)$
gives $R\leq(1-R)/3+o(1)$, which is asymptotically $R\leq 1/4$, essentially equivalent to [^3] Strictly speaking, to obtain the positive $\epsilon$ in (8) we need the fact that the Plotkin bound is not tight at positive rates. (8). The stronger bound (9) is obtained instead by using the best known bound on $\delta$, which is the linear programming bound of [26] adapted to $q$-ary codes [21], defined implicitly in $\delta_{2}$ by the inequality $R\leq R_{LP1}(q,\delta_{2})$ with $q=3$ and
@@
$H_{q}(t)=t\log_{q}(q-1)-t\log_{q}t-(1-t)\log_{q}(1-t)\,.$
The bound on $R$ is found by combining (18) with $R\leq R_{LP}(3,\delta_{2})$, which gives $R\leq\frac{1}{2}\delta^{*}$, where $\delta^{*}$ is the solution of the equation $\frac{1}{2}\delta=R_{LP1}(3,\delta)$.
@@
The usefulness of this approach can be appreciated from the possible extensions that it allows. The general case is presented in the next Section, but we anticipate here the result obtained for ternary linear codes with asymptotic $3$-hash relative distance $\delta_{3}$. To do this, we need the following generalization of Jamison’s bound to multiple coverings, due to Bruen [27].
Lemma 2 ([27])
Let $\mathcal{H}$ be a multiset of hyperplanes in ${\mathbb{F}}_{q}^{m}$. If no hyperplane in $\mathcal{H}$ contains 0 and each point in ${\mathbb{F}}_{q}^{m}\setminus\{0\}$ is covered by at least $t>0$ hyperplanes in $\mathcal{H}$, then
$|\mathcal{H}|\geq(m+t-1)(q-1)\,.$
Theorem 2
For $\delta_{3}>0$, the rate of ternary codes with asymptotic $3$-hash relative distance $\delta_{3}$ is bounded above as $R\leq\frac{1}{2}\delta^{*}-\delta_{3}$, where $\delta^{*}$ is the unique solution of the equation
$\frac{1}{2}\delta-\delta_{3}=R_{LP1}(3,\delta)\,.$
Proof:
The proof is a simple extension of the one presented to rederive (9). Assume that the code $C$ has $3$-hash distance at least $d_{3}$. Then each codeword of the subcode $C^{\prime}$ (defined as before equation (17)) must contain now at least $d_{3}$ symbols equal to $2$ in the first $d_{2}$ coordinates, that is, each nonzero element of $\mathbb{F}_{3}^{m-1}$ is covered at least $d_{3}$ times by the hyperplanes in (17). Using Lemma 2 we deduce now $d_{2}\geq 2(m+d_{3}-2)$, which in terms of asymptotic rates with $d_{3}/n\to\delta_{3}$ gives R≤12δ_2-δ_3 +o(1) . Then, combining this bound with the linear programming bound $R\leq R_{LP1}(q,\delta_{2})$ leads to the statement. ∎
This bound improves the one derived in [18] for linear codes, but is not tight as $\delta_{3}$ approaches $2/9$, becoming worse than the bounds in [7] and [20], which hold even for general codes, see Figure 1.
@@ III General Case $q\geq k\geq 3$
We now turn to the case of $q$-ary linear codes proving our main result, which is a relation between the rate and minimum $k$-hash distances of codes for any $q\geq k\geq 3$. This will then be instantiated both to the case of asymptotic relative distance $\delta_{k}>0$ and to the case where only the requirement $d_{k}\geq 1$ is considered. Our results will follow from the following basic lemma.
@@
Let $0,x_{1},x_{2},\ldots,x_{s-1}$ be $s$ codewords with $s$-hash distance $d_{s}>0$. Assume without loss of generality that they are all pairwise distinct in the first $d_{s}$ coordinates. Let $C^{\prime}$ be a subcode of dimension (ms+1)(m-s+1) which only intersects trivially (in the origin) the subspace spanned by $x_{1},x_{2},\ldots,x_{s-1}$. Define, for $i=1,\ldots,d_{s}$, the set $S_{i}={\mathbb{F}}_{q}\setminus\{0,x_{1,i},x_{2,i},\ldots,x_{s-1,i}\}$. Let $G$ be the (full rank) $(m-s+1)\times n$ generator matrix of the subcode $C^{\prime}$ and $g_{i}$ be its $i$-th column. Consider the hyperplanes H_i,b={v∈F_q^m-s+1 ∣v⋅g_i = b} , i=1,…,d_s , b∈S_i . Since $|S_{i}|=q-s$, these are $(q-s)d_{s}$ hyperplanes, none of which contains the zero vector. Assume $d_{s+1}>0$; then, each $v\in{\mathbb{F}}_{q}^{m-s+1}\setminus\{0\}$ is covered at least $d_{s+1}$ times by those hyperplanes. Therefore, from Lemma 2 we have (q-s)d_s≥(m-s+1+d_s+1-1)(q-1) . This is equivalent to d_s+1 ≤q-sq-1d_s-m+s , whenever the right hand side is positive. Otherwise, the assumption $d_{s+1}>0$ is impossible, and hence $d_{s+1}=0$. ∎
Note that Lemma 3 is an improvement on the recursive bound given in equation (12) whenever $d_{s}\geq q-1$. Iteration of Lemma 3 gives the following result, from which all our bounds will be derived.
@@
For a $q$-ary linear code of dimension $m$ with minimum Hamming distance $d_{2}$, the $k$-hash distance satisfies the inequality
$d_{k}\leq\frac{(q-2)^{\underline{k-2}}}{(q-1)^{k-2}}\left(d_{2}-\sum_{i=1}^{k-2}(m-i-1)\frac{(q-1)^{i}}{(q-2)^{\underline{i}}}\right)^{+}\,.$ (20)
Note that Theorem 3 is an improvement of (13). Whenever we assume $m/n\to R$ and $d_{k}/n\to\delta_{k}>0$ as $n\to\infty$, after some manipulations equation (20) gives the following result.
@@
Consider linear codes in ${\mathbb{F}}_{q}^{n}$ of rate $R$ and asymptotic minimum relative $k$-hash distance $\delta_{k}>0$ and asymptotic minimum Hamming relative distance $\delta_{2}>0$. Then, as $n\to\infty$,
$R\leq\frac{1}{\sum_{i=1}^{k-2}\frac{(q-1)^{i}}{(q-2)^{\underline{i}}}}\left(\delta_{2}-\frac{(q-1)^{k-2}}{(q-2)^{\underline{k-2}}}\delta_{k}\right)+O\left(\frac{1}{n}\right)\,.$
For example, when applied for $q=7$ and $k=4$, the above equation says
$R\leq\frac{1}{3}\delta_{2}-\frac{3}{5}\delta_{4}+o(1)\,$,
while equation (13) gives
$R\leq\frac{1}{2}\delta_{2}-\frac{1}{2}\delta_{4}+o(1)\,.$
A comparison of the $(R,\delta_{4})$–bounds derived when combining these two equations with the bound $R\leq R_{LP1}(7,\delta_{2})$ is shown in Figure 2.
Figure 2: Bounds on the rate of linear $7$-ary codes with $4$-hash distance $\delta_{4}$.
If we instead simply assume $d_{k}\geq 1$ as $n\to\infty$, Theorem 3 gives the following result (which matches the expression obtained in the limit $\delta_{k}\to 0$ in Corollary 1).
Corollary 2
Consider linear $k$-hash codes in ${\mathbb{F}}_{q}^{n}$ of rate $R=m/n$ and relative minimum Hamming distance $\delta_{2}$. Then, as $n\to\infty$,
$R\leq\frac{\delta_{2}}{\sum_{i=1}^{k-2}\frac{(q-1)^{i}}{(q-2)^{\underline{i}}}}+O\left(\frac{1}{n}\right)\,.$
It can be easily seen that this bound improves the bound valid for general codes of equation (14) for every $q$ and $k$.
In the rest of this section, we study the upper bounds on the rate of linear $k$-hash codes that result from Corollary 2 once we combine it with different bounds on the minimum Hamming distance. It is worth pointing out, indeed, that different such bounds lead to the stronger conclusions in different regimes. Using the Plotkin bound for $q\geq 3$, for example, we obtain the following corollary.
@@
Now, using the upper bound on $\delta$ of equation (22) and the bound for linear $k$-hash codes given in Corollary 2 we obtain R ≤q-1q (1-R)∑i=1k-2(q-1)i(q-2)i + o(1) . Rearranging the terms we obtain the statement of the corollary. ∎
@@
where $\delta^{*}$ is the unique root of the equation in $\delta$
$\delta=\left(\sum_{i=1}^{k-2}\frac{(q-1)^{i}}{(q-2)^{\underline{i}}}\right)R_{LP1}(q,\delta)$
where $0\leq\delta\leq\frac{q-1}{q}$ and $R_{LP1}$ is the function defined in equation (19).
Remark 1
We observe that one could use the second linear programming bound or the straight-line bounds given in [21, 28] to improve the results of Corollaries 3 and 4 for different values of $q$ and $k$. Here we refrain from showing those improvements to keep a simpler presentation of our results.
In Table I, we compare the bounds provided in Corollaries 34 and the one given in (4) for $k=3$ and $q\in[3,64]$. It can be seen that the linear programming bound performs better for $q\leq 19$ while for $q\geq 23$ the Plotkin bound gives a better result.
@@
33 $1/4=0.25$ 0.2198 0.3691
44 $1/3\approx 0.3334$ 0.3000 $1/2=0.5$
55 $3/8=0.375$ 0.3441 0.5694
77 $5/12\approx 0.4167$ 0.3928 0.6438
88 $3/7\approx 0.4286$ 0.4080 $2/3\approx 0.6667$
99 $7/16=0.4375$ 0.4200 0.6846
@@
$13$ $11/24\approx 0.4584$ 0.4497 0.7298
$16$ $7/15\approx 0.4667$ 0.4628 $3/4=0.75$
$17$ $15/32\approx 0.4688$ 0.4663 0.7554
1919 $17/36\approx 0.4723$ 0.4721 0.7646
$23$ $21/44\approx 0.4773$ 0.4811 0.7790
$25$ $23/48\approx 0.4792$ 0.4846 0.7847
$27$ $25/52\approx 0.4808$ 0.4877 0.7897
$29$ $27/56\approx 0.4822$ 0.4903 0.7942
$31$ $29/60\approx 0.4834$ 0.4927 0.7982
$32$ $15/31\approx 0.4839$ 0.4938 $4/5=0.8$
$37$ $35/72\approx 0.4862$ 0.4984 0.8081
$41$ $39/80\approx 0.4875$ 0.5013 0.8134
@@
$64$ $31/63\approx 0.4921$ 0.5119 $5/6\approx 0.8334$
Figure 3: Bounds on the rate of $q$-ary linear $4$-hash codes for $q\geq 5$.
For a fixed value of $k$ and $q\to\infty$, both bounds given in Corollaries 3 and 4 converge to 1/(k1)1/(k-1), which is the same upper bound on the rate of $(q,k)$-hash codes (not necessarily linear) that one can derive from equation (5). Corollary 3 approaches 1/(k1)1/(k-1) from below since the rhs of equation (21) is strictly increasing in $q$ and since $\sum_{i=1}^{k-2}\frac{(q-1)^{i}}{(q-2)^{\underline{i}}}\geq k-2$, while Corollary 4 does not, see for example Table I where for $k=3$ and $q=41,\ldots,64$ we have upper bounds that exceed $1/2$.
@@
since the function $\log_{x}(\alpha x)$ is increasing in $x$ for $x\geq 2$ and $0<\alpha<1$ and since $k\geq 4$ and $q\geq k^{2}$. Then, by Bernoulli’s inequality we have that 12 ( q-k+2q )^k-2 ≥12(1 - (k-2)2q) . Since the left hand side of (23) is less than 1/(k1)1/(k-1), in order to prove the statement of the theorem we just need to show that 1k-1 ≤12(1 - (k-2)2q) , but this inequality is satisfied for $q\geq k^{2}$} and $k\geq 4$. ∎
@@
We conjecture that Theorem 4 still holds also if we relax the hypothesis to $q\geq 2k-3$ and $k\geq 3$. This would imply that, for all the interesting values of $q$ and $k$ (since the asymptotic rate of linear $k$-hash codes in ${\mathbb{F}}_{q}^{n}$ for $q\leq 2k-4$ is zero), our bound provides the best result. In support of our conjecture, Table I provides an instance for $k=3$ where for every $q\geq 3$ the bound of Corollary 3 improves the one of equation (4) and Figure 3 reports the comparison between our bounds and the best known bounds in the literature for $k=4$ and $q\geq 2k-3$. In addition, we have numerically verified the conjecture for every $k\in[3,100]$ and $q\geq 2k-3$. In Figure 3 , we compare our upper bounds for $k=4$ and $q\geq 2k-3$. We note that for a fixed value of $k$ both the upper bounds of Corollaries 3, 4 and the lower bounds derived in [7, Th. 2] approach 1/(k1)1/(k-1) as $q\to\infty$.
IV A Related Open Problem in Zero-Error Capacity
As mentioned in the introduction, the problem of perfect hashing can be alternatively described as a problem of zero-error capacity under list decoding for a certain class of discrete memoryless channels. Consider for example two discrete memoryless channels as in Figure 4, where arrows represent transitions with positive probability. For both channels, any two inputs symbols are “confusable”, in the sense that they share at least one output. This means that any two input sequences of length $n$ are also confusable, in the sense that there is an output sequence which is compatible with both. If the decoder is required to output one single estimated input sequence, with a positive probability there will be an error. In list-decoding with list-size $L$, instead, the decoder is allowed to output a list of $L$ possible input sequences, and an error is declared if the true input sequence is not in the list. A zero-error code is one for which the decoder is always able to produce a list of $L$ codewords which include the one sent by the encoder. This can be equivalently stated by saying that for any output sequence there are at most $L$ compatible input sequences. For the channel on the left of Figure 4 with list-size $L=2$, we see that zero-error codes are precisely the trifferent codes discussed in Section II-B. Indeed, a code is zero-error with $L=2$ if no output sequence is compatible with more than two distinct codewords, which means that for any three codewords there is a coordinate where they all differ, making one of them incompatible with the observed output. A similar reasoning shows that also for the channel on the right, zero-error codes are possible with list-size $L=2$, and these are precisely the $(4,3)$-hash codes. More generally, $(q,k)$-hash codes are zero-error codes with decoding list-size k1k-1 for symmetric channels with $q$ inputs and $\binom{q}{k-1}$ outputs, each one reachable from k1k-1 inputs.